One of the air-purification methods for enclosed spaces involves the use of "scrubbers" containing aqueous lithium hydroxide, which reacts with carbon dioxide to produce lithium carbonate and water: 2 LiOH(aq) + 2 CO2(g) → Li2CO3(s) + H2O(l) Consider the air supply in a submarine with a total volume of 2.20 × 105 L. The pressure is 0.9970 atm, and the temperature is 25° C. If the pressure in the submarine drops by 0.0100 atm as the result of carbon dioxide being consumed by an aqueous lithium hydroxide scrubber, how many moles of CO2 are consumed?_______________mol
Answer #1

step 1: calculate initial mol of CO2

Given:

P = 0.997 atm

V = 220000 L

T = 25.0 oC

= (25.0+273) K

= 298 K

find number of moles using:

P * V = n*R*T

0.997 atm * 220000 L = n * 0.08206 atm.L/mol.K * 298 K

n = 8.97*10^3 mol

step 2: calculate final mol of CO2

Given:

P = 0.987 atm

V = 220000.0 L

T = 25.0 oC

= (25.0+273) K

= 298 K

find number of moles using:

P * V = n*R*T

0.987 atm * 220000 L = n * 0.08206 atm.L/mol.K * 298 K

n = 8.88*10^3 mol

So,

mol of CO2 removed = final mol - initial mol

= 8.97*10^3 mol - 8.88*10^3 mol

= 90.0 mol

Answer: 90.0 mol